OpenStudy (anonymous):

Find the value of y for the given value x Y=x+9, x=23

OpenStudy (uri):

y=23+9 y=32

HanAkoSolo (jamierox4ev3r):

just plug in the x-variable y=23+9 y=32

HanAkoSolo (jamierox4ev3r):



HanAkoSolo (jamierox4ev3r):

does this make sense @flower93 ?

OpenStudy (anonymous):

Awesome, Thanks you!

OpenStudy (uri):

Are you saying i didn't make sense? :(

OpenStudy (anonymous):

THank you for your help @uri :)) sorry